4998 472.000 Operating income 1 000 Opersting inceme Amount of Change
Answer #1

Amount of change (in millions)

(recent year -prior year)

% of change increase or decrease
Revenue 499488-473000=26488 26488/472000=.056 or 5.6% increase
Operating expense 421296-402000=19296 19296/402000=.048 or 4.8‰ increase
Operating income 78192-71000=7192 7192/71000= .101 or 10.1% Increase

B)

During the recent year, revenue has increased and operating expense has increased.

As a Result ,operating income has increased. The increasing trend from the prior year.